# Dichotomy theorem distinguishing non-integrability and the lowest-order Yang-Baxter equation for isotropic spin chains

### v1

@@ Abstract
We investigate the integrability and non-integrability of isotropic spin chains with nearest-neighbor interaction with general spin $S$. We prove a dichotomy theorem that a single relation sharply separates two scenarios: (i) this system is non-integrable, or (ii) the lowest-order Yang-Baxter equation is satisfied. This result solves in the affirmative the Grabowski-Mathieu conjecture stating that a model is integrable only if this model has a 3-local conserved quantity. This theorem also serves as a complete classification of integrability and non-integrability for $S\leq 13.5$, suggesting that all the integrable models are in the scope of the Yang-Baxter equation.
@@
Introduction.— Integrability and non-integrability of quantum many-body systems are important subjects in theoretical physics. Integrable systems have been studied for almost 100 years, and various quantum systems such as the Heisenberg model, the transverse Ising model, and more complicated models were discovered as integrable models JM ; Bax . We note that unlike classical integrability there is no universally-accepted definition of quantum integrability and non-integrability, though known integrable models with local interactions share common properties including solvability and infinitely many local conserved quantities. Following previous studies CM11 ; GE16 ; Shi19 , in this Letter, we adopt a working definition of integrability and non-integrability as having infinitely many and no nontrivial local conserved quantities, respectively.
@@
To find and solve integrable models, a sophisticated technique based on the Yang-Baxter equation, named the algebraic Bethe ansatz or the quantum inverse scattering methods, works as a powerful tool Fad ; KBI . Precisely, if the Yang-Baxter equation is satisfied, then we can obtain an infinite series of local conserved quantities DG ; FF ; GM94 ; GM95-1 ; NF20 , with which energy eigenstates can be constructed. In this case, we also have a boost operator, which generates local conserved quantities through a simple calculation GM94 ; GM95-1 . It is noteworthy, however, that some important integrable models including the Hubbard model and the XY model neither satisfy the conventional (differentiable) Yang-Baxter equation nor have a boost operator GM95-2 ; Lin01 ; Lee21 .
In contrast to integrability, theoretical aspects of non-integrability have not yet been fully addressed, though non-integrability (absence of nontrivial local conserved quantities) is a relevant property to various phenomena including thermalization Lan ; EF ; SM17 ; MS17 ; SMreply and response to external drives Kub ; SK ; SF ; Zot . One reason for this insufficiency is that the proof of non-existence (no solution or no nontrivial local conserved quantity) is usually harder than the proof of existence (of solutions and nontrivial local conserved quantities). Despite the scarcity of studies focusing on non-integrability, the existing literature can broadly be classified into two approaches.
The first approach to non-integrability is a traditional one, examining solvability and unsolvability by restricting the solution method to the Yang-Baxter equation KS82 ; Ken92 ; BY94 ; BN14 ; Hie92 ; LPR19 ; Lee20 ; Lee21 ; Mai24 ; LP24 . With this approach, several classes of spin chains have been classified into solvable and unsolvable ones in terms of the Yang-Baxter equation. As far as previous research has explored, all models satisfying the lowest-order Yang-Baxter equation (the Reshetikin condition KS82 ; Ken92 ) also satisfy the full-order Yang-Baxter equation, implying the solvability. However, this approach suffers from an inherent limitation: we cannot rule out the possibility that a model is not solvable by the Yang-Baxter equation but is still solvable by another, possibly yet unknown, solution method.
The second approach possesses generality, where the presence or absence of local conserved quantities is directly examined. A pioneering investigation in this direction was carried out by Grabowski and Mathieu GM95-2 , where they put two conjectures on spin chains with nearest-neighbor interaction: (1) A model is integrable only if it has a 3-local conserved quantity. (2) A model is integrable if it satisfies the lowest-order Yang-Baxter equation. Conjecture 1 directly concerns non-integrability, while conjecture 2 is closely aligned with the field of integrable systems, which is related to the first approach. Henceforth, we refer to conjecture 1 as the Grabowski-Mathieu conjecture otherwise explicitly noted. The Grabowski-Mathieu conjecture and its extension GP21 , if proved, are expected to show its significance when it is applied to previously unexplored systems, since it serves as the guidepost of investigation.
Recently, a mathematical technique to prove non-integrability was invented Shi19 , and by applying this technique various quantum spin systems have been proven to be non-integrable Chi24-1 ; Shi24 ; PL24-1 ; CY24 ; ST24 ; Chi24-2 ; PL24-2 ; HYC24 ; YCS24-1 ; YCS24-2 ; Shi25 ; FT25 . Notably, symmetric $S=1/2$ systems in one YCS24-1 ; YCS24-2 and higher dimensions ST24 ; Chi24-2 are completely classified into integrable and non-integrable, which confirms the Grabowski-Mathieu conjecture in the affirmative within this class. Another important progress is brought by Hokkyo Hok25 , which observes similarity among many previous non-integrability proofs and clarifies the sufficient conditions for proving non-integrability. He showed that if two operator equations respectively have solutions only in specific forms, then the same non-integrability proof works. However, these classification theorems are derived by exhaustively proving the integrability and non-integrability of all models in this class, which differs from the original aim of the Grabowski-Mathieu conjecture as the guidepost for investigation of wide classes of systems that include many unexplored systems. An exceptional general statement Hok25 employs not simple criteria, and thus what concrete systems adopt these results are still not clear.
@@
In this Letter, we bridge the existing two approaches and establish a dichotomy theorem for isotropic spin chains with general spin $S$. Precisely, we prove that whether a single equation contains a certain type of term or not completely determines whether this model is non-integrable or satisfies the lowest-order Yang-Baxter equation. As its direct consequence, any integrable system is shown to have 3-local and 4-local conserved quantities generated by a boost operator, and hence the Grabowski-Mathieu conjecture is solved in the affirmative for isotropic spin chains. Importantly, the separation into the two extreme scenarios applies all isotropic spin chains including unexplored ones, which indeed meets the aim of the Grabowski-Mathieu conjecture. We also note that our criterion is tested by a simple and direct calculation, which shows clear contrast to existing results GM95-2 ; Hok25 requiring to solve search problems of operator equations.
@@
Figure 1: Claim of our dichotomy theorem. If $D$ has no 3-support operator, then this model satisfies the lowest-order yang-Baxter equation (the Reshetikin condition) and has a boost operator generating 3-local and 4-local conserved quantities. If $D$ has a 3-support operator, then this model is non-integrable in the sense that this model has no nontrivial local conserved quantity.
@@
Consider a shift-invariant isotropic spin chain with general spin $S$ with nearest-neighbor interaction, whose Hamiltonian is given by Eq. (1). For the above class of spin chains, whether the quantity
$D=[[h_{i,i+1},h_{i+1,i+2}],h_{i,i+1}+h_{i+1,i+2}]$ (2)
contains a nonzero 3-support operator or not completely determines whether the model is non-integrable or satisfies the lowest-order Yang-Baxter equation (the Reshetikin condition).
The latter case is equivalent to the existence of a boost operator in the form of $B=\sum_{j}jh_{j,j+1}$ generating 3-local and 4-local conserved quantities given by
@@
$Q^{4}=\sum_{i}$ $2[[h_{i,i+1},h_{i+1,i+2}],h_{i+2,i+3}]$
$+[[h_{i,i+1},h_{i+1,i+2}],h_{i+1,i+2}].$ (4)
This boost operator is conjectured to generate higher conserved quantities $Q^{n}$ for $n\geq 5$ (conjecture 2 of Ref. GM95-2 ), while whether the generated operators are indeed conserved quantities is an open problem. The details of the lowest-order Yang-Baxter equation (the Reshetikin condition) and the boost operator are presented in SM .
This theorem reveals that the calculation of a single relation completely determines either non-integrability or the lowest-order Yang-Baxter equation. This shows a high contrast to the Grabowski-Mathieu conjecture GM95-2 and its partial proof by Hokkyo Hok25 , where we need to solve a search problem on operator equations. In addition, provided conjecture 2 of Grabowski and Mathieu GM95-2 , our theorem implies an astonishing claim that whether $D$ contains a nonzero 3-support operator or not completely determines the integrability and non-integrability of isotropic spin chains.
We here briefly comment on what happens if the above conjecture is false in isotropic spin chains. In this case, there exists an isotropic spin chain with a finite number of nontrivial local conserved quantities, which is called a partially-integrable system YCS24-1 ; YCS24-2 . Moreover, this counterexample is higher than or equal to $S=14$. These observations impress that conjecture 2 is highly plausible.
Implications of this theorem.— This theorem has various implications for integrability and non-integrability.
First, this theorem confirms that all integrable isotropic spin chains should have 3-local and 4-local conserved quantities in the form of Eqs. (3) and (4). Our theorem excludes the existence of (unknown) integrable systems without 3-local and 4-local conserved quantities. This proves the Grabowski-Mathieu conjecture GM95-2 in the affirmative in the class of isotropic spin chains, stating that a system is integrable only if it has a 3-local conserved quantity. This is the first proof of the Grabowski-Mathieu conjecture for a large class of spin systems including models which we have not yet explored.
In addition, our theorem fills the gap between the two approaches to non-integrability explained in the introduction. In particular, this dichotomy theorem justifies the validity of the first approach with fixing solution method Ken92 ; BY94 by confirming that the lowest-order Yang-Baxter equation does not miss integrable systems in isotropic spin chains. Given the fact that some integrable models including the Hubbard model do not satisfy the lowest-order Yang-Baxter equation, it is striking that all isotropic integrable models admit the lowest-order Yang-Baxter equation.
This theorem also serves as the complete classification of integrability and non-integrability up to $S=13.5$. Kennedy Ken92 and Batchelor and Yung BY94 have examined all isotropic spin chains satisfying the Reshetikin condition less than or equal to $S=13.5$ by computer search and found that all the spin chains satisfying the Reshetikin condition are in fact exhausted by known integrable systems solvable by the Yang-Baxter equation. Precisely, there are four infinite sequences of integrable systems and one exceptional integrable system for $S=3$ (see End matter), which correspond to four infinite series of compact Lie groups and one exceptional Lie group satisfying SU(2) symmetry BY94 . Combining this fact and our theorem, we conclude that these four infinite sequences of integrable systems and one exceptional integrable system are the complete list of integrable isotropic spin chains up to $S=13.5$.
Reshetikin condition and the Yang-Baxter equation.— Before going to the proof, we here clarify the relationship among conditions on integrable systems. The Yang-Baxter equation is a relation on a one-parameter family of a matrix $R_{ij}(\lambda)$ on two sites $i$ and $j$ given by
@@
where $h$ is the local Hamiltonian and $\Pi$ is a swap operator. The third and higher-order terms are not universal.
Plugging Eq. (6) into the Yang-Baxter equation (5) and compare the lowest nontrivial order, $\lambda\mu^{2}$, we obtain the Reshetikin condition claiming that $D$ in Eq. (2) can be expressed as KS82 ; Ken92
$D_{123}=E_{23}-E_{12},$ (7)
where $E_{ij}$ is a proper operator on two sites $i$ and $j$. Thus, the Reshetikin condition is the nontrivial lowest-order relation of the Yang-Baxter equation. The Reshetikin condition is also equivalent to the relation $[Q^{3},H]=0$ with $Q^{3}$ given by Eq. (3), meaning the presence of 3-local conserved quantity in this form. Note that the Reshetikin condition also implies $[Q^{4},H]=0$ with $Q^{4}$ given by Eq. (4) (see Supplemental Material SM ).
@@
Proof: former part.— We first consider the case that $D$ does not have 3-support operators with nonzero coefficients. By comparing our condition to the lowest-order Yang-Baxter equation (7), our task boils down to that 2-support operators in $D$ take the form of $E_{i+1,i+2}-E_{i,i+1}$. We shall show this fact as follows.
@@
which follows the argument below. Since both $[h_{i,i+1},h_{i+1,i+2}]$ and $h_{i+1,i+2}$ are $SU(2)$ symmetric, their commutator $[[h_{i,i+1},h_{i+1,i+2}],h_{i+1,i+2}]$ should also be $SU(2)$ symmetric. Notice that all 2-support $SU(2)$ symmetric operator takes the form of the right-hand side of Eq. (1), which also has inversion symmetry.
Proof: latter part.— We next consider the case that $D$ has a 3-support operator with nonzero coefficients. We prove the non-integrability of this Hamiltonian with the help of the scheme proposed by Hokkyo Hok25 . Below we denote the $m$-support part of $k$-local conserved quantity by $Q^{k,m}$, i.e., we expand $Q^{k}$ as
@@
We first restrict the possible form of $k$-support operator $Q^{k,k}$. We here introduce the doubling-product operators $\mathit{\Delta}^{k}$ which is recursively defined as
@@
It is shown in Ref. Hok25 that $Q^{k,k}$ should be proportional to $\mathit{\Delta}^{k}$ if $f(A)=[I_{1}\otimes A_{2},h_{12}]$ is injective and that the equation for 2-support operator $X$
@@
Discussion.— We have proven that a single quantity $D$ given in Eq. (2) sharply distinguishes two possible scenarios of isotropic spin chains: If $D$ has a 3-support operator, then the spin chain is non-integrable, i.e., no nontrivial local conserved quantity. If $D$ has no 3-support operator, then the spin chain satisfies the lowest-order Yang-Baxter equation (the Reshetikin condition) and has 3-local and 4-local conserved quantities in the form of Eqs. (3) and (4). Our result solves the Grabowski-Mathieu conjecture GM95-2 in the affirmative for isotropic spin chains, which is the first proof of the Grabowski-Mathieu conjecture applicable to systems whose integrability has still been unknown. Our result also serves as the classification theorem of integrability and non-integrability for the case with less than or equal to $S=13.5$.
In our theorem, a single relation serves as a criterion to distinguish the lowest-order Yang-Baxter equation and non-integrability, which is an advantage compared to previously proposed criteria GM95-2 ; Hok25 which contain search problems of operator equations. Provided conjecture 2 by Grabowski and Mathieu GM95-2 , the integrability or non-integrability of isotropic spin chains is completely determined by the same single relation. Such a simple criterion of integrability and non-integrability for a wider class of spin systems beyond isotropic spin chains is worth investigating.
Our result shows that all integrable isotropic spin chains satisfy the lowest-order Yang-Baxter equation. Moreover, restricting isotropic spin chains with $S\leq 13.5$, all integrable isotropic spin chains satisfy the Yang-Baxter equation. This shows high contrast to the Hubbard model, which is solvable but does neither satisfy the lowest-order Yang-Baxter equation nor have a boost operator generating local conserved quantities. On the other hand, the Hubbard model admits a generalized version of the Yang-Baxter equation and a generalized boost operator Lin01 ; Lee21 . On the basis of the above, our theorem based on the non-integrability proof sheds new light on the strongness and the limitation of the Yang-Baxter equation from the perspective of non-integrability, implying that the Yang-Baxter equation is the only solution method of integrable systems in some classes. We note that the technique of non-integrability proofs is also extended to write down local conserved quantities of integrable systems NF20 ; YF23 . The non-integrability proof approach will pave the way to clarify how universally the Yang-Baxter equation applies, which merits further research.
@@
“Dichotomy theorem distinguishing non-integrability and the lowest-order Yang-Baxter equation for isotropic spin chains”

### v2

# Dichotomy theorem separating complete integrability and non-integrability of isotropic spin chains

@@ Abstract
We investigate the integrability and non-integrability of isotropic spin chains with nearest-neighbor interaction with general spin $S$ in terms of the presence or absence of local conserved quantities. We prove a dichotomy theorem that whether a single quantity is zero or not sharply separates two scenarios: (i) this system has $k$-local conserved quantities for all $k$ (completely integrable), or (ii) this system has no nontrivial local conserved quantity (non-integrable). This result excludes the possibility of an intermediate system with some but not all local conserved quantities, which solves in the affirmative the Grabowski-Mathieu conjecture. This theorem also serves as a complete classification of integrability and non-integrability for $S\leq 13.5$, suggesting that all the integrable models are in the scope of the Yang-Baxter equation.
@@
Introduction.— Integrability and non-integrability of quantum many-body systems are important subjects in theoretical physics. Integrable systems have been studied for almost 100 years, and various quantum systems such as the Heisenberg model, the transverse Ising model, and more complicated models were discovered as integrable models JM ; Bax . We note that unlike classical integrability there is no universally-accepted definition of quantum integrability and non-integrability, though known integrable models with local interactions share common properties including solvability and infinitely many local conserved quantities. Following previous studies CM11 ; GE16 ; Shi19 , in this Letter, we characterize complete integrability and non-integrability as having all $k$-local conserved quantities and no nontrivial local conserved quantities, respectively.
@@
To find integrable models, a sophisticated technique based on the Yang-Baxter equation, named the algebraic Bethe ansatz or the quantum inverse scattering methods, works as a powerful tool Fad ; KBI . Precisely, if the Yang-Baxter equation is satisfied, then we can construct an infinite series of local conserved quantities DG ; FF ; GM94 ; GM95-1 ; NF20 , with which energy eigenstates can also be constructed. In this case, we also have a boost operator, which generates local conserved quantities through a simple calculation GM94 ; GM95-1 . It is noteworthy, however, that some important integrable models including the Hubbard model and the XY model also have an infinite family of local conserved quantities but neither satisfy the conventional (differentiable) Yang-Baxter equation nor have a boost operator GM95-2 ; Lin01 ; Lee21 .
In contrast to integrability, theoretical aspects of non-integrability have not yet been fully addressed, though the absence of nontrivial local conserved quantities is a relevant property to various phenomena including thermalization Lan ; EF ; SM17 ; MS17 ; SMreply and response to external drives Kub ; SK ; SF ; Zot . Within the sparse literature devoted to non-integrability, one traditional approach is restricting the solution method to the Yang-Baxter equation and testing the validity of its lowest-order relation (the Reshetikin condition) KS82 ; Ken92 ; BY94 ; BN14 ; Hie92 ; LPR19 ; Lee20 ; Lee21 ; Mai24 ; LP24 . However, this approach suffers from an inherent limitation: we cannot rule out the possibility that a model is not solvable by the Yang-Baxter equation but is still solvable by another, possibly yet unknown, solution method. A more recent approach directly proves the absence of nontrivial local conserved quantities, which was invented by Ref Shi19 , and by applying this technique various quantum spin systems have been proven to be non-integrable Chi24-1 ; Shi24 ; PL24-1 ; CY24 ; ST24 ; Chi24-2 ; PL24-2 ; HYC24 ; YCS24-1 ; YCS24-2 ; Shi25 ; FT25 . Notably, symmetric $S=1/2$ systems in one YCS24-1 ; YCS24-2 ; Shi25 and higher dimensions ST24 ; Chi24-2 are completely classified into completely integrable and non-integrable. Another important progress is brought by Hokkyo Hok25 , which observes similarity among many previous non-integrability proofs and clarifies the sufficient conditions for proving non-integrability. Precisely, it is shown that there are two key operator equations and if these two operator equations respectively have specific forms of solutions as unique solutions, then the non-integrability proof surely works.
In principle, there may exist an intermediate system which has some $k$-local conserved quantities but not all ones. On the other hand, all known models have (i) all $k$-local conserved quantities, or (ii) no nontrivial local conserved quantity, and we have never found such intermediate systems. Therefore, we strongly expect the dichotomy that only the above two extremes are possible. A similar conjecture was carried out by Grabowski and Mathieu GM95-2 , observing that the presence or absence of a 3-local conserved quantity determines the complete integrability or non-integability. However, proving this dichotomy on a rigorous foundation has been left as an open problem.
@@
In this Letter, we rigorously establish this dichotomy for isotropic spin chains with general spin $S$. Precisely, we prove that whether a single quantity $D^{3}$ is zero or not fully determines whether this model has all $k$-local conserved quantities (completely integrable) or no nontrivial local conserved quantity (non-integrable). This confirms the all-or-nothing principle of local conserved quantities, excluding the possibility of an intermediate system with some local conserved quantities. The separation into the two extreme scenarios applies all isotropic spin chains including unexplored ones, which indeed meets the aim of the Grabowski-Mathieu conjecture. Notably, our criterion of this dichotomy is easy to compute, which shows clear contrast to existing attempts GM95-2 ; Hok25 requiring to solve complex search problems of operator equations.
@@
We here introduce a key quantity
$D=[[h_{i,i+1},h_{i+1,i+2}],h_{i,i+1}+h_{i+1,i+2}]$ (2)
and denote its 3-support operator terms by $D^{3}$.
Figure 1: Claim of our dichotomy theorem. If $D^{3}=0$, then this model has all $k$-local conserved quantities (completely integrable). If $D^{3}\neq 0$, then this model has no nontrivial local conserved quantity (non-integrable).
@@
Consider a shift-invariant isotropic spin chain with general spin $S$ with nearest-neighbor interaction, whose Hamiltonian is given by Eq. (1). For the above class of spin chains,
  • If $D^{3}=0$, then this model has $k$-local conserved quantities for all positive integers $k\geq 3$ (completely integrable)
  • If $D^{3}\neq 0$, then this model has no nontrivial local conserved quantity (non-integrable)
Here, in the latter case we exclude all $k$-local conserved quantities with $13\leq k\leq N/2-$ 1. Our result establishes the dichotomy of local conserved quantities (all or nothing), excluding a partially-integrable system YCS24-1 ; YCS24-2 with a finite number of nontrivial local conserved quantities. In addition, the criterion of this dichotomy is whether a single quantity is zero or not, which is easy to calculate. This shows a high contrast to the Grabowski-Mathieu conjecture GM95-2 ; GP21 and its partial proof by Hokkyo Hok25 , where we need to solve search problems on operator equations to confirm non-integrability.
Note that in the case of $D^{3}=0$ the $k$-local conserved quantities $Q^{k}$ is generated recursively by the boost operator in the form of $B=\sum_{j}jh_{j,j+1}$ as
$Q^{k+1}=[B,Q^{k}]$ (3)
with $Q^{2}=H$. The 3-local conserved quantity, for example, is given by
@@
with which $[Q^{3},H]$ is equivalent to the shift sum of $D$ given in Eq. (2).
Yang-Baxter equation and Reshetikin condition.— In order to explain the significance of our result, we here briefly introduce the Yang-Baxter equation and the Reshetikin condition (for details, see the Supplemental Material SM ). The Yang-Baxter equation is a relation on a one-parameter family of a matrix $R_{ij}(\lambda)$ on two sites $i$ and $j$ given by
@@
where $h$ is the local Hamiltonian and $\Pi$ is a swap operator. The third and higher-order terms are not universal. Plugging Eq. (6) into the Yang-Baxter equation (5) and compare the lowest nontrivial order, $\lambda\mu^{2}$, we obtain the Reshetikin condition claiming that $D$ in Eq. (2) can be expressed as KS82 ; Ken92
$D_{i,i+1,i+2}=E_{i+1,i+2}-E_{i,i+1},$ (7)
where $E_{ij}$ is a proper operator on two sites $i$ and $j$.
Implications of the main theorem.— Remarkably, our theorem confirms that the Reshetikin condition does not miss integrable systems in isotropic spin chains, justifying the approach to non-integrability based on the Reshetikin condition explained in the introduction (i.e., asserting non-integrability by the violation of the Reshetikin condition) Ken92 ; BY94 . The integrable part ($D^{3}=0$) of our theorem contains the Reshetikin condition, since the Reshetikin condition (7) is equivalent to the condition $[Q^{3},H]=0$ with Eq. (4). Note that our dichotomy theorem imposes that if an infinite family of local conserved quantities exists, the Reshetikin condition should be satisfied. Hence, our theorem establishes that the Reshetikin condition is the necessary and sufficient condition for complete integrability in isotropic spin chains. In light of the fact that some integrable models including the Hubbard model do not satisfy the Reshetikin condition, it is striking that all isotropic integrable models admit the Reshetikin condition.
In addition, our theorem proves the Grabowski-Mathieu conjecture (both conjecture 1 and 2) GM95-2 ; GP21 in the affirmative in the class of isotropic spin chains, stating that a system is integrable if and only if it has a 3-local conserved quantity. This is the first proof of the Grabowski-Mathieu conjecture for a large class of spin systems including models which we have not yet explored.
This theorem also serves as the complete classification of complete integrability and non-integrability up to $S=13.5$. Kennedy Ken92 and Batchelor and Yung BY94 have examined all isotropic spin chains satisfying the Reshetikin condition less than or equal to $S=13.5$ by computer search and found that all the spin chains satisfying the Reshetikin condition are in fact exhausted by known integrable systems solvable by the Yang-Baxter equation. Precisely, there are four infinite sequences of integrable systems and one exceptional integrable system for $S=3$ (see End matter), which correspond to four infinite series of compact Lie groups and one exceptional Lie group satisfying SU(2) symmetry BY94 . Combining this fact and our theorem, we conclude that these four infinite sequences of integrable systems and one exceptional integrable system are indeed the complete list of integrable isotropic spin chains up to $S=13.5$.
@@
Proof for $D^{3}=0$.— We here consider the case of $D^{3}=0$. We here employ a recent result Hok25-2 stating that if the 3-local conserved quantity $Q^{3}$ given by Eq. (4) conserves ($[Q^{3},H]=0$), then all $k$-local conserved quantities $Q^{k}$ given by Eq. (3) also conserve ($[Q^{k},H]=0$) (see Lemma. S.1 in the Supplemental Material SM ). Hence, it suffices to show the presence of a 3-local conserved quantity, or the Reshetikin condition. By comparing the condition $D^{3}=0$ to the Reshetikin condition (7), our task boils down to that 2-support operators in $D$ take the form of $E_{i+1,i+2}-E_{i,i+1}$. We shall show this fact as follows.
@@
under which we can set $A_{ij}=E_{ij}$. We can show the inversion symmetry as follows. Since both $[h_{i,i+1},h_{i+1,i+2}]$ and $h_{i+1,i+2}$ are $SU(2)$ symmetric, their commutator $[[h_{i,i+1},h_{i+1,i+2}],h_{i+1,i+2}]$ should also be $SU(2)$ symmetric. Notice that all 2-support $SU(2)$ symmetric operator takes the form of the right-hand side of Eq. (1), which also has inversion symmetry.
Proof for $D^{3}\neq 0$.— We next consider the case of $D^{3}\neq 0$. We prove the non-integrability of this Hamiltonian with the help of the scheme proposed by Hokkyo Hok25 . Below we denote the $m$-support part of $k$-local conserved quantity by $Q^{k,m}$, i.e., we expand $Q^{k}$ as
@@
We first restrict the possible form of the $k$-support operator $Q^{k,k}$. We here introduce the doubling-product operators $\mathit{\Delta}^{k}$ which is recursively defined as
@@
It is shown in Ref. Hok25 that $Q^{k,k}$ should be proportional to $\mathit{\Delta}^{k}$ if $f(A)=[I_{1}\otimes A_{2},h_{12}]$ is injective and the equation for 2-support operator $X$
@@
Discussion.— We have proven that an isotropic spin chain has (i) all $k$-local conserved quantities (completely integrable), or (ii) no nontrivial local conserved quantity (non-integrable). We have also shown that whether a single quantity $D^{3}$ given in Eq. (2) is zero or not sharply distinguishes these two possible scenarios. Our result solves the Grabowski-Mathieu conjecture GM95-2 ; GP21 in the affirmative for isotropic spin chains. Our result also serves as the classification theorem of integrability and non-integrability for the case with less than or equal to $S=13.5$.
In our theorem, a single quantity $D^{3}$ serves as a criterion to distinguish the complete integrability and non-integrability, which is an advantage compared to previously proposed criteria GM95-2 ; GP21 ; Hok25 containing search problems of operator equations. One may expect an extension of the same criterion to a wider class of systems beyond isotropic ones. However, the Hubbard model, which has many symmetries close to isotropy, has all $k$-local conserved quantity but $D^{3}\neq 0$. This reflects the fact that the local conserved quantities of the Hubbard model are generated not by the standard boost operator $B$ but its generalized one Lin01 ; Lee21 . In spite of this negative suggestion, it is still plausible that a simple criterion similar to $D^{3}$ determines integrability and non-integrability for a wider class of spin systems beyond isotropic spin chains, which is worth investigating.
Our result shows that all integrable isotropic spin chains satisfy the lowest-order Yang-Baxter equation (the Reshetikin condition). Moreover, restricting isotropic spin chains with $S\leq 13.5$, all integrable isotropic spin chains satisfy the Yang-Baxter equation, meaning that the Yang-Baxter equation is the unique solution method in this regime. On the other hand, again the Hubbard model does not admit the original Yang-Baxter equation Lee21 . On the basis of the above, we expect that the Yang-Baxter equation is not always universal but indeed universal in some classes including isotropic spin chains. We note that the technique of non-integrability proofs is also extended to write down local conserved quantities of integrable systems NF20 ; YF23 . The non-integrability proof approach will pave the way to clarify how universally the Yang-Baxter equation applies, which merits further research.
@@
“Dichotomy theorem separating complete integrability and non-integrability of isotropic spin chains”